OpenStudy (anonymous):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128 Show work

OpenStudy (anonymous):

first of what is the common ratio?

OpenStudy (anonymous):

by what number do you multiply to obtain the next term? (this number is called a 'common ratio' and denoted with r)

OpenStudy (anonymous):

the first term, denoted as \(a_1\) and in this case, as you can see, it is -2.

OpenStudy (anonymous):

So, all you need to do, is to fill the form below, for the first term and common ratio \[\large a_n=a_{{\tiny~}_1} \left({\rm r} \right)^{n-1}\]

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

@Nnesha can you help please

Nnesha (nnesha):

okay first you need to find ( ratio) geometric sequence mean when you get next term by dividing or multiplying previous term by something so first term is -2 what you multiply with -2 to get next term which is 8 ?

OpenStudy (anonymous):

i have no actual clue on how to do these. soooooo im guessing the next step would be like 8x4????

OpenStudy (anonymous):

sorry to bother u but i have alot of questions i need help on today and your like the only person i know on here that actually helps me figure stuff out..

Nnesha (nnesha):

-2 times what = 8 8 times what = -32

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

-2x-2=8 8x-4=32 -32x4=128 ???

OpenStudy (anonymous):

i think i missed a - sign or two

Nnesha (nnesha):

yep right -4 \[\huge\rm 2 \times -4 =8\] \[\huge\rm 8 \times -4 =-32\] \[\huge\rm -32 \times -4 = 128\]

Nnesha (nnesha):

now plug r value into the formula that @((*-*)) gave you

OpenStudy (anonymous):

r value?

Nnesha (nnesha):

\[\huge\rm a_n = a_1 \times r^{n-1}\] r = ration a_1 = is first term

Nnesha (nnesha):

r = ratio

OpenStudy (anonymous):

now i feel dumb because i have no idea what a ratio is

Nnesha (nnesha):

ratio:t he common ratio is the ratio of a term to the previous term http://www.regentsprep.org/regents/math/algtrig/atp2/geoseq.htm

OpenStudy (anonymous):

does that mean the ratio is -2 and the common ratio is -4?

OpenStudy (anonymous):

an= -2 x (-4) ^(n-1) is this correct?

Nnesha (nnesha):

yep right

Nnesha (nnesha):

no commone ratio is r and -2 is first term

OpenStudy (anonymous):

can you show me exactly how to write this on my home work and show all my work correctly

Nnesha (nnesha):

there is not a lot work to who i already showed that how to find common ratio then you just have to use formula to plug in a_1 and r values

Nnesha (nnesha):

show*

OpenStudy (anonymous):

so i show my work like -2x-2=8 8x-4=-32 -32x-4=128 an=-2x(-4)^(n-1) or do i just write an=-2x(-4)^(n-1)

Nnesha (nnesha):

yep right looks good

OpenStudy (anonymous):

thanks. is it okay to tag you in a few other questions later on? it should be just check my work questions.

Nnesha (nnesha):

sure :-)